# Is attention all you need to solve the correlated electron problem?

### v1

@@ I Introduction
Recently, neural network (NN) based variational Monte Carlo (VMC) has been developed for solving the many-electron problem with high accuracy [4, 5, 6, 7, 8, 9, 10, 11]. Compared to human-designed trial wavefunctions, neural network wavefunctions contain a large number of parameters, has enormous representation power, and can be optimized efficiently. NN-VMC has been shown to be highly accurate in calculating the ground state energy of interacting electrons in atoms and molecules [9, 10, 11, 12, 13], lattice models [6, 14, 15], uniform electron gas [16, 17, 18, 19, 20, 21], moire semiconductors [22, 23], and fractional quantum Hall liquids [24, 25].
@@
The performance of our attention based NN ansatz is evaluated for interacting electrons in semiconductor moiré heterobilayers, such as WSe2/WS2. This moiré platform hosts a fascinating variety of correlated electronic states, including Mott insulators [27, 28], generalized Wigner crystals [29], and strongly correlated Fermi liquids [30]. Here, doped electrons reside on one semiconductor layer and experience a moiré potential. The Hamiltonian thus takes the form of a two-dimensional Coulomb electron gas in a periodic potential [31]:
@@ III Neural Network Wavefunction
Figure 1: Architecture of the neural network wavefunction ansatz. SlaterNet: Multilayer perceptron neural network generates one-body orbitals to approximate general single Slater determinant wavefunctions. Psi-Solid: Self-attention neural network for solids based on Psi-Former [11] capturing correlations by mixing individual particle streams.
@@
Recently, a new type of variational method based on neural network wavefunctions has been developed to approximate the ground state of many-electron systems. Traditional variational wavefunctions for correlated systems are artfully designed by humans, tailor made for a given quantum phase, and usually contain only a few variational parameters. In contrast, thanks to their universal approximation capability, deep neural networks can be used to generate powerful wavefunction ansatz, which has a large number of variational parameters and can accurately represent distinct phases of matter in a unified way [34, 10].
@@ III.1 Deep neural network for one-body orbitals
for any particle $i$, where the two vectors $\bm{L}=n\bm{L}_{1}+m\bm{L}_{2}$ with $n,m\in\mathbb{Z}$ specify the supercell size and geometry. For periodic systems, we choose the function ${\rm feature}(\bm{r})$ to express the coordinate $\bm{r}$ in terms of periodic coordinates [35],
@@ III.3 Self-attention neural network for electron correlations
${\cal N}=\sqrt{d_{\rm head}}\sum_{j=1}^{N}\exp({\bm{q}_{j}^{lh}\cdot\bm{k}_{i}^{lh}})\,.$
@@
where the accumulation is performed by the learned transformation $W_{\rm o}^{l}\in\mathbb{R}^{d_{\rm L}}\times\mathbb{R}^{N_{\rm heads}d_{\rm head}}$ projecting the output of the attention heads onto the dimension of $\bm{h}_{i}^{l}$. Equations (19) and (20) replace the compositional relation Eq. (13) in the construction of one-body orbitals as in SlaterNet. By mixing different single-particle streams with the self-attention mechanism, correlations between the particle states are captured.
@@
for any $k,l\neq i$. The permutation equivariance is necessary for the generalized Slater determinant to represent a fermionic wavefunction that is antisymmetric under permutation of particle coordinates. This is property is inherent to the functional form of the self-attention operation Eq. (18).
@@ V Results
This section presents the numerical results obtained from our variational Monte Carlo calculations of Hartree-Fock using SlaterNet and the minimal self-attention NN Psi-Solid and the benchmark with band-projected exact diagonalization.
@@ V.1 Convergence and scaling with system size
Figure 3: (a) Raw (unsmoothed) learning curves for the energy per electron at $\nu=2/3$ filling for 6 electrons with $\epsilon=5$, and 18 electrons with $\epsilon=10$. (b) Converged energy of the moiré system at $\nu=2/3$ filling and $\epsilon=10$ as a function of the number of variational parameters in Psi-Solid. Error bars indicate the standard deviation of the mean local energy, averaged over the batch size at the end of optimization. Stars mark the estimated number of parameters, $N^{*}_{\text{par}}$, required to reach convergence saturation. Further increasing in the number of parameters keeps the energies within one standard deviation (highlighted by the striped region). For sake of presentation, the $8e$ data was shifted up by $0.39$meV, and $18e$ data shifted down by $0.16$meV.
@@
To further validate our approach, we examine the convergence of the ground-state energy per particle as a function of the number of variational parameters for the 9-, 12-, and 27-site systems at $\nu=2/3$ filling. Fig. 3(b) explores networks spanning three orders of magnitude in variational parameters and reveals a clear convergence pattern in all cases, with the ground-state energy approaching a lower bound as the network reaches 1 million variational parameters.
@@
$N_{\text{par}}^{*}\approx 320\times N^{2.1}\ .$ (32)
@@
While this scaling law provides a useful estimate of the number of parameters required to achieve a given accuracy, the mere parameter count is not the only relevant factor. Since attention heads play a crucial role in capturing correlations, a minimum number of attention heads and layers (approximately 2–3 in our tests) is necessary to reach the ground state. Simply increasing the total number of parameters without ensuring sufficient attention heads and layers would not improve performance. Our scaling law serves as a guideline within the regime where these minimal requirements are met. We leave a more detailed analysis of scaling laws as a function of individual network dimensions for future work.
Figure 4: Scaling law for the number of variational parameters at convergence saturation as a function of the number of electrons. The extracted saturation points are well described by a power-law relation indicating how the required parameter count grows with system size. This scaling provides a guideline for selecting an optimal network size while ensuring convergence
@@ V.2 Benchmark for small system size
We study a variety of different system sizes and regimes at different interaction strengths. The comprehensive energy comparisons given in Fig. LABEL:fig:Method_Comparison graphically evince the difference in ground state energy levels between all 3 methods for 9 unit cell system again at $\nu=2/3$. The self-attention NN energy is considerably lower than scHF, with the difference precisely representing the correlation energy. We see that the correlation energy is about $2\%$ in our moiré system, considerably higher than the correlation energy in ordinary molecular systems [3]. This enhanced correlation energy underscores the relevance of moiré systems for studying strongly correlated phenomena. Additionally, we observe that as the number of bands increases in BP-ED, the ground-state energy approaches but remains higher than that of the self-attention NN .
@@
Figure 5: Comparison of ground state energies obtained from self-attention neural network Psi-Solid (blue), Hartree-Fock from SlaterNet (green), and BP-ED (dots) in a $3\times 3$ supercell with $\nu=2/3$ filling. Dashed lines for Psi-Solid and SlaterNet indicate mean of local energy obtained from the variational Monte Carlo algorithm. The shaded region indicates the standard deviation of the mean of the local energy averaged over the batch size per optimization step. Dashed gray line is a guide to the eyes.
$N_{e}$ $\epsilon$  Psi-Solid  SlaterNet  BP-ED
@@ V.3 Fermi Liquids and Generalized Wigner Crystals
Finally, the difference in energies from the self-attention NN and scHF allows to estimate correlation energy Eq. (3). We include the extracted correlation energy as a function of dielectric constant in App. D.
@@ VI Comparison to other neural network ansätze
In contrast to previous works studying periodic solids with neural quantum states [35, 39], we do not multiply an envelope function when constructing single-particle orbitals. Instead the full functional dependence of the periodic orbitals is represented by the neural network. Ref. [35] motivated the use of envelope functions by a better convergence for their system; however our numerical experiments for our system did not show any improvements when envelope functions are included. Envelope functions introduce human bias towards a certain functional dependence, therefore we remove them in our network architecture.
@@ Appendix B Jastrow factors
In systems with periodic boundary condition, the electron-electron distance $r$ that enters the Jastrow factor is calculated with respect to a periodic and smooth norm that reproduces the Euclidean norm in the limit $r\to 0$. The smooth and periodic norm $||\bm{r}||_{M}$ with respect to the lattice vectors $M=(\bm{m}_{1},\bm{m}_{2})$ is defined as [35]
@@
where $G=2\pi M^{-1}$ are the reciprocal lattice vectors.
@@ Appendix C Hyperparameters
of Psi-Solid Attention heads per layer $6$
@@
Table 2: Table of default hyperparameters used in our numerical calculations.
Table 2 presents the default hyperparameters used in our calculations, ensuring stable and monotonous convergence during training. As discussed in the main text, architectural parameters must be adjusted according to system size to achieve convergence. For the convergence and scaling law analysis, we explored various hyperparameter distributions. The table lists the set that successfully ensured convergence for the largest system studied, consisting of 18 electrons.
@@ Appendix D Correlation Energy
Here, we investigate the correlation energy per electron, defined as the ground-state energy difference between the Hartree-Fock (SlaterNet) and self-attention NN (Psi-Solid) approaches. We analyze its dependence on the interaction strength, controlled by the dielectric constant $\epsilon$, see Fig. 8 for the result. As $\epsilon$ decreases (corresponding to stronger interactions), the correlation energy gradually increases, reflecting the growing importance of many-body effects beyond mean-field theory. This trend highlights the limitations of Hartree-Fock in capturing strong correlations and underscores the role of self-attention NN in accurately describing the correlated electronic states.
@@
Figure 8: Correlation energy per electron, $E_{\text{corr}}=E_{\text{Psi-Solid}}-E_{\text{SlaterNet}}$, at $\nu=2/3$ filling with $27$ sites as a function of inverse dielectric constant $\epsilon^{-1}$.

### v2

@@ I Introduction
Recently, neural network (NN) based variational Monte Carlo (VMC) has been developed for solving the many-electron problem with high accuracy [4, 5, 6, 7, 8, 9, 10, 11, 12]. Compared to human-designed trial wavefunctions, neural network wavefunctions contain a large number of parameters, has enormous representation power, and can be optimized efficiently. NN-VMC has been shown to be highly accurate in calculating the ground state energy of interacting electrons in atoms and molecules [9, 10, 11, 13, 14], lattice models [6, 15, 12, 16, 17], uniform electron gas [18, 19, 20, 21, 22, 23, 24], moire semiconductors [25, 26], and fractional quantum Hall liquids [27, 28].
@@
The performance of our attention based NN ansatz is evaluated for interacting electrons in semiconductor moiré heterobilayers, such as WSe2/WS2. This moiré platform hosts a fascinating variety of correlated electronic states, including Mott insulators [30, 31, 32], generalized Wigner crystals [33], and strongly correlated Fermi liquids [34]. Here, doped electrons reside on one semiconductor layer and experience a moiré potential. The Hamiltonian thus takes the form of a two-dimensional Coulomb electron gas in a periodic potential [35]:
@@ III Neural Network Wavefunction
Figure 1: Architecture of the neural network wavefunction ansatz. SlaterNet: Multilayer perceptron neural network generates one-body orbitals to approximate general single Slater determinant wavefunctions. Self-attention NN: Self-attention neural network for solids based on Psiformer [11] capturing correlations by mixing individual particle streams.
@@
Recently, a new type of variational method based on neural network wavefunctions has been developed to approximate the ground state of many-electron systems. Traditional variational wavefunctions for correlated systems are artfully designed by humans, tailor made for a given quantum phase, and usually contain only a few variational parameters. In contrast, thanks to their universal approximation capability, deep neural networks can be used to generate powerful wavefunction ansatz, which has a large number of variational parameters and can accurately represent distinct phases of matter in a unified way [38, 10, 39].
@@ III.1 Deep neural network for one-body orbitals
for any particle $i$, where the two vectors $\bm{L}=n\bm{L}_{1}+m\bm{L}_{2}$ with $n,m\in\mathbb{Z}$ specify the supercell size and geometry. For periodic systems, we choose the function ${\rm feature}(\bm{r})$ to express the coordinate $\bm{r}$ in terms of periodic coordinates [pescia2022neural, 40],
@@ III.3 Self-attention neural network for electron correlations
${\cal N}=\sqrt{d_{\rm AttnValues}}\sum_{j=1}^{N}\exp({\bm{q}_{j}^{lh}\cdot\bm{k}_{i}^{lh}})\,.$
@@
where the accumulation is performed by the learned transformation $W_{\rm o}^{l}\in\mathbb{R}^{d_{\rm L}}\times\mathbb{R}^{N_{\rm heads}d_{\rm AttnValues}}$ projecting the output of the attention heads onto the dimension of $\bm{h}_{i}^{l}$. Equations (19) and (20) replace the compositional relation Eq. (13) in the construction of one-body orbitals as in SlaterNet. Optionally, multiple perceptron layers can be applied between the attention layers, which is realized by multiple subsequent applications of (20) with individually learnable parameters. By mixing different single-particle streams with the self-attention mechanism, correlations between the particle states are captured.
@@
for any $k,l\neq i$. The permutation equivariance is necessary for the generalized Slater determinant to represent a fermionic wavefunction that is antisymmetric under permutation of particle coordinates. This property is inherent to the functional form of the self-attention operation Eq. (18).
@@ V Results
This section presents the numerical results obtained from our variational Monte Carlo calculations of Hartree-Fock using SlaterNet and the minimal self-attention NN and the benchmark with band-projected exact diagonalization.
@@ V.1 Convergence and scaling with system size
Figure 3: (a) Raw (unsmoothed) learning curves for the energy per electron at $\nu=2/3$ filling for 6 electrons with $\epsilon=5$, and 18 electrons with $\epsilon=10$. (b) Converged energy of the moiré system at $\nu=2/3$ filling and $\epsilon=10$ as a function of the number of variational parameters in the self-attention NN. Error bars indicate the standard deviation of the mean local energy, averaged over the batch size at the end of optimization. Stars mark the estimated number of parameters, $N^{*}_{\text{par}}$, required to reach convergence saturation. Further increasing in the number of parameters keeps the energies within one standard deviation (highlighted by the striped region). For sake of presentation, the $8e$ data was shifted up by $0.39$meV, and $18e$ and $24e$ datas shifted down by $0.16$meV and $0.25$meV, respectively. Default simulation parameters are contained in Table 2.
@@
To further validate our approach, we examine the convergence of the ground-state energy per particle as a function of the number of variational parameters for the 9-, 12-, 27- and 36-site systems at $\nu=2/3$ filling. Fig. 3(b) explores networks spanning three orders of magnitude in variational parameters and reveals a clear convergence pattern in all cases, with the ground-state energy approaching a lower bound as the network reaches 1 million variational parameters.
@@
$N_{\text{par}}^{*}\approx 400\times N^{2.0}\ .$ (32)
@@
While this scaling law provides a useful estimate of the number of parameters required to achieve a given accuracy, the mere parameter count is not the only relevant factor. Since attention heads play a crucial role in capturing correlations, a minimum number of attention heads and layers (approximately above 3 in our tests) is necessary to reach the ground state. Simply increasing the total number of parameters without ensuring sufficient attention heads and layers would not improve performance. Our scaling law serves as a guideline within the regime where these minimal requirements are met. We leave a more detailed analysis of scaling laws as a function of individual network dimensions for future work.
Figure 4: Scaling law for the number of variational parameters at convergence saturation as a function of the number of electrons for $\epsilon=10$. The extracted saturation points are well described by a power-law relation indicating how the required parameter count grows with system size. This scaling provides a guideline for selecting an optimal network size while ensuring convergence.
@@ V.2 Benchmark for small system size
We study a variety of different system sizes and regimes at different interaction strengths. The comprehensive energy comparisons given in Fig. 5 graphically evince the difference in ground state energy levels between all 3 methods for 9 unit cell system again at $\nu=2/3$. The self-attention NN energy is considerably lower than scHF, with the difference precisely representing the correlation energy. We see that the correlation energy is about $2\%$ in our moiré system, considerably higher than the correlation energy in ordinary molecular systems [3]. This enhanced correlation energy underscores the relevance of moiré systems for studying strongly correlated phenomena. Additionally, we observe that as the number of bands increases in BP-ED, the ground-state energy approaches but remains higher than that of the self-attention NN .
@@
Figure 5: Comparison of ground state energies obtained from the self-attention NN (blue), Hartree-Fock from SlaterNet (green), and BP-ED (dots) in a $3\times 3$ supercell with $\nu=2/3$ filling. Dashed lines for the self-attention NN and SlaterNet indicate mean of local energy obtained from the variational Monte Carlo algorithm. The shaded region indicates the standard deviation of the mean of the local energy averaged over the batch size per optimization step. Dashed gray line is a guide to the eyes.
$N_{e}$ $\epsilon$  self-attention NN  SlaterNet  BP-ED
@@ V.3 Fermi Liquids and Generalized Wigner Crystals
Finally, the difference in energies from the self-attention NN and scHF allows to estimate correlation energy Eq. (3). We include the extracted correlation energy as a function of dielectric constant in App. D. Furthermore, in App. E, we provide additional exact-diagonalization results on the metal-insulator phase transition.
@@ VI Comparison to other neural network ansätze
In contrast to previous works studying periodic solids with neural quantum states [40, 44], we do not multiply an envelope function when constructing single-particle orbitals. Instead the full functional dependence of the periodic orbitals is represented by the neural network. Ref. [40] motivated the use of envelope functions by a better convergence for their system; however our numerical experiments for our system did not show any improvements when envelope functions are included. Ref. [44] found faster convergence using plane wave envelopes when the system was close to a Hartree-Fock state. Envelope functions introduce human bias towards a certain functional dependence, therefore we remove them in our network architecture.
@@ Appendix B Jastrow factors
In systems with periodic boundary condition, the electron-electron distance $r$ that enters the Jastrow factor is calculated with respect to a periodic and smooth norm that reproduces the Euclidean norm in the limit $r\to 0$. The smooth and periodic norm $||\bm{r}||_{M}$ with respect to the supercell vectors $M=(\bm{L}_{1},\bm{L}_{2})$ is defined as [40]
@@
where $G=2\pi M^{-1}$ are the reciprocal supercell vectors.
@@ Appendix C Hyperparameters
Attention heads per layer $6$
@@
Table 2: Table of default hyperparameters used in our numerical calculations with the self-attention neural network.
Table 2 presents the default hyperparameters used in our calculations, ensuring stable and monotonous convergence during training. As discussed in the main text, architectural parameters must be adjusted according to system size to achieve convergence. For the convergence and scaling law analysis, we explored various hyperparameter distributions. The table lists the set that successfully ensured convergence for the 18 electron system. These parameters are applied to generate the figures in the main text.
@@
The Fig. 8 contains the precise network dimensions used to generate the parameter scaling in Fig. 3 in the main text.
Figure 8: Converged energy of the moiré system at $\nu=2/3$ filling and $\epsilon=10$ as a function of number of variational parameters. The labels indicate the parameters $(n_{\text{layers}},n_{\text{heads}},\text{dim}_{\text{att}},\text{dim}_{\text{perc}})$ used for each calculation. The total number of parameters are counted explicitly from the checkpoints saved during the training.
@@ Appendix D Correlation Energy
Here, we investigate the correlation energy per electron, defined as the ground-state energy difference between the Hartree-Fock (SlaterNet) and self-attention NN approaches. We analyze its dependence on the interaction strength, controlled by the dielectric constant $\epsilon$, see Fig. 9 for the result. As $\epsilon$ decreases (corresponding to stronger interactions), the correlation energy gradually increases, reflecting the growing importance of many-body effects beyond mean-field theory. This trend highlights the limitations of Hartree-Fock in capturing strong correlations and underscores the role of self-attention NN in accurately describing the correlated electronic states.
@@
Figure 9: Correlation energy per electron, $E_{\text{corr}}=E_{\text{self-attn NN}}-E_{\text{SlaterNet}}$, at $\nu=2/3$ filling with $27$ sites as a function of inverse dielectric constant $\epsilon^{-1}$.
Appendix E BP-ED Benchmark
In addition to benchmarking our self-attention NN energies with BP-ED, we leverage BP-ED to reveal deeper details on our system, namely identifying the Fermi-Liquid (FL) and Generalized Wigner Crystal (GWC) phases, along with the order and location of the metal-insulator (MIT) transition. Fig. 10 shows two distinct phase signatures at different interaction strengths $\epsilon=10$ [(a),(c)], and $\epsilon=5$ [(b),(d)]. By assigning each crystal momenta an integer index, the Many Body Spectra (MBS) in Fig. 10 ([a),(b)] shows the lowest energies per momentum sector, with the corresponding momentum space positions of each sector graphically displayed in the moiré Brillouin zone (MBZ) as a function of ground state (GS) occupation in Fig. 10 [(c),(d)].
The 6-fold rotational symmetry of the 27 site cluster allows the GS to accommodate a 6-fold degeneracy, emblematic of a high mobility, Fermi-Liquid-like state. On the other hand, the 27 site cluster also accommodates a tripled unit cell where the corners of the MBZ $\kappa,\kappa^{\prime}$ fold back to $\gamma$ in the BZ, indicating an expected 3-fold GS degeneracy at $\gamma,\kappa,\kappa^{\prime}$ for a GWC signature. Not only does the MBS in Fig 10 [(a),(b)] show clear signatures of a FL and GWC state at respective interactions, but also Fig 10 (c) shows the presence of a Fermi Surface while its $\epsilon^{-1}=5$ counterpart does not.
The exact MIT point predicted by BP-ED is shown in Fig. 11, in which the many-body energy gap (difference between 4th and 3rd lowest energy values) are plotted versus interaction strength $\epsilon^{-1}$. Our expectation of an MIT as interaction strength increases is validated by the 6-fold GS degeneracy eventually giving way to a 3-fold GS degeneracy at about $\epsilon^{-1}=0.11$, signaling the crossover from FL to GWC. Our BP-ED results are well consistent with our self-attention NN results in Fig. 6, as evidence for FL and GWC states are demonstrated at the appropriate interaction strengths, showcasing the ability of our self-attention NN to capture the physics of this moiré system.
Figure 10: Many Body Spectrum (MBS) [(a),(b)] and ground state occupation [(c),(d)] for 27 sites at 2/3 filling for (above) $\epsilon=10$ and (below) $\epsilon=5$. The MBS of the left figure demonstrates the 6-fold GS degeneracy, while the MBS on the right demonstrates the 3-fold GS degeneracy at the high symmetry points, indicating a Generalized Wigner Crystal state. A clear Fermi Surface is also visible in (c) as the system is in a Fermi Liquid state.
Figure 11: Many Body Energy Gap ($E_{gap}=E_{4}-E_{3}$, where $E_{i}$ labels the state with the $i^{th}$ lowest energy) as a function of interaction strength $\epsilon^{-1}$. The gap opens just after $\epsilon^{-1}=0.10$, signaling an MIT at about $\epsilon^{-1}=0.11$.
BP-ED techniques can also be used to gain insight on the nature of this MIT. We can characterize the transition as first or second order by probing the discontinuity of observables across the transition point. First order transitions are characterized by discontinuities across transition points while second order transitions are continuous across transition points for large enough system size. Since we are dealing with finite system sizes in BP-ED, there will always be a discontinuity; however if this discontinuity grows with system size, then we can deduce that correlation length will not tend to 0 with increasing system size, thereby showing evidence for a first order transition. Inversely, a decrease in the observable discontinuity across the transition point as system size grows is evidence for a second order phase transition.
By probing the kinetic plus periodic potential one-body energy ($H_{0}$ in equation 1):
$\left\langle H_{0}\right\rangle=\sum_{\mathbf{k}}\epsilon_{\mathbf{k}}\left\langle c_{\mathbf{k}}^{\dagger}c_{\mathbf{k}}\right\rangle_{\mathbf{GS}},$ (49)
and the band projected structure factor:
$\bar{S}(\mathbf{q})=\frac{1}{N_{e}}\left\langle\bar{n}(\mathbf{q})\bar{n}(\mathbf{-q})\right\rangle_{\mathbf{GS}},$ (50)
across the MIT point ($\epsilon^{-1}=0.11$) for 21 and 27 site clusters (21, 27-k-point mesh detailed in [61]), we can measure the discontinuity gap to shed light on the order of the phase transition. Since the projected structure factor has peaks along $\mathbf{q}=\kappa,\kappa^{\prime}$, we plot $\bar{S}(\mathbf{q}=\kappa)$ as a function of interaction strength to isolate the effect of the transition.
FIG. 12 not only demonstrates a discontinuity exactly at the critical point for each cluster, but also shows that the discontinuity gap grows as system size grows for both kinetic energy and the band projected structure factor (for $\mathbf{q}=\kappa$). The discontinuity gap for $\langle H_{0}\rangle$ increases from 0.61 meV to 0.87 meV and from 0.88 to 1.35 for $\bar{S}(\mathbf{q}=\kappa)$, demonstrating evidence for a first-order MIT.
Figure 12: $\langle H_{0}\rangle$ [(a),(c)] and $S(\mathbf{q}=\kappa)$ [(b),(d)] as a function of $\epsilon^{-1}$ for the 21 [(a-b)] and 27 site [(c-d)] systems. Here, $\langle...\rangle_{GS}$ denotes the expectation value over a single ground state. The discontinuity gap increases with system size evincing a first order metal-insulator transition.